OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (1, -5)

OpenStudy (ddcamp):

Horizontal lines have a slope of 0. So, using point slope form: (y-y₁) = m(x-x₁) y-(-5) = 0(x-1) y+5 = 0 y = -5

OpenStudy (anonymous):

horizontal line passing through the point (1, -5)... horizontal line means y = something the y value of the point is -5, that's the something thus y = -5.

OpenStudy (anonymous):

something has to be a constant